In Kylie's house, the hot water is set at 130 degrees. The incoming cold water is 40 degrees. How much hot and cold water is needed to make a 10 gallon bath at 108 degrees?
What I've tried:
h= # of gallons of hot water c= # of gallons of cold water
130h+40c=108(10)

Answer by richwmiller(17219)

That is a good start.
h+c=10,
130h+40c=108(10)
c = 22/9, h = 68/9
c=2.4444, h=7.5556